an equipotential surface that surrounds a +3.0"uC" point charge has a radius of 2.0cm. what is the potential of this surface?


1
Expert's answer
2021-04-28T13:04:41-0400
"V=k\frac{q}{ r}\\V=9\cdot10^9\frac{3\cdot10^{-6}}{ 0.02}=1.35\cdot10^6\ V"